OpenStudy (anonymous):

Find a third–degree polynomial function such that f(0) = –12 and whose zeros are 1, 2, and 3. Using complete sentences, explain how you found it.

OpenStudy (anonymous):

f(x)=k(x-1)(x-2)(x-3) f(0)=k(-1)(-2)(-3)=-12 k=2 f(x)=2(x-1)(x-2)(x-3)

OpenStudy (anonymous):

Thank you so much